OpenStudy (anonymous):

simplify 6 radical 27

OpenStudy (precal):

\[6\sqrt{27}\]

OpenStudy (precal):

\[6\sqrt{3*3*3}\]

OpenStudy (precal):

\[6*3\sqrt{3}\]

OpenStudy (precal):

\[18\sqrt{3}\]